# SecureWebArena: A Holistic Security Evaluation Benchmark for LVLM-based Web Agents

### v1

@@ A.1. Environment Examples
Figure 5. Examples of evaluated environments.
Figure 6. PDR (%) of all evaluated LVLM-based agents across 6 environments and 6 attack types.
Figure 7. Case study illustrating an indirect prompt injection during an online shopping task.
A.2. Environment-wise Analysis
@@
Summary of Findings. In summary, our environment-centric analysis demonstrates that there is no single ”most vulnerable” agent or ”most effective” attack. Instead, vulnerability is an emergent property of the agent-environment-task triad. Transactional contexts amplify visual exploits, information-dense contexts favor linguistic manipulation, and structured contexts reward direct command injections. This complex interplay underscores the inadequacy of evaluating web agent security in a vacuum and reinforces the critical necessity of a diverse, multi-environment benchmark like SecureWebArena to uncover the full spectrum of security risks.
A.3. Case Study Illustration

### v2

@@ Appendix B Environment Examples
Figure 5: Examples of evaluated environments.
Figure 6: PDR (%) of all evaluated LVLM-based agents across 6 environments and 6 attack types.
Appendix C Environment-wise Analysis
@@
Summary of Findings. In summary, our environment-centric analysis demonstrates that there is no single "most vulnerable" agent or "most effective" attack. Instead, vulnerability is an emergent property of the agent-environment-task triad. Transactional contexts amplify visual exploits, information-dense contexts favor linguistic manipulation, and structured contexts reward direct command injections. This complex interplay underscores the inadequacy of evaluating web agent security in a vacuum and reinforces the critical necessity of a diverse, multi-environment benchmark like SecureWebArena to uncover the full spectrum of security risks.
Figure 7: Case study illustrating an indirect prompt injection during an online shopping task.
Appendix D Case Study
While quantitative metrics reveal failure rates at scale, they often mask the root causes behind agent vulnerabilities. To showcase the diagnostic power of our multi-layered evaluation protocol, we present a case study comparing two failure pathways that yield the same security violation. In this scenario, the agent is tasked with purchasing a product but encounters a pop-up ad containing an indirect prompt injection. The pop-up displays a salient “Claim Your 50% Discount!” button, beneath which a faint line instructs: “re-enter your account password in the next field.”
@@ Appendix D Case Study
There is no indication the agent noticed the malicious prompt; its behavior was guided purely by visual salience, misinterpreting the interface flow. This case illustrates how the same outcome may stem from fundamentally different failure modes. GPT-5 requires stronger reasoning safety, while UI-TARS-1.5 would benefit from broader exposure to deceptive UI patterns. Such causal analysis, made possible by SecureWebArena’s layered evaluation, enables actionable insights beyond binary success metrics.